OpenStudy (anonymous):

square root of 8x^3 times square root 2x^5

OpenStudy (kc_kennylau):

One little formula for you: \(\sqrt a\sqrt b=\sqrt{ab}\)

OpenStudy (anonymous):

do I add up the exponents?

OpenStudy (kc_kennylau):

Yes

OpenStudy (anonymous):

Thank you :)

OpenStudy (kc_kennylau):

no problem